OpenStudy (anonymous):

find the inverse of the function f(x)= 2+√x

OpenStudy (mathmate):

y=(x-2)^2

OpenStudy (anonymous):

oh thank you! :P

OpenStudy (anonymous):

are there any other steps before getting to the result?

OpenStudy (mathmate):

y=2+sqrt(x) switch x and y x=2+sqrt(y) solve for y sqrt(y)=x-2 y=(x-2)^2

OpenStudy (anonymous):

thanx man you saved me!

OpenStudy (mathmate):

You're welcome!

OpenStudy (anonymous):

man if you could check out the other questions i have it would be great. Thank you again :D

OpenStudy (nikvist):

\[f^{-1}(x)=(x-2)^2\quad,\quad x\geq 2\]